OpenStudy (anonymous):

The expression radical 50 + radical 32 is equivalent to what?

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{50}\] simplifies to what?

OpenStudy (anonymous):

\[5\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

what about \[\large \sqrt{32}\]

OpenStudy (anonymous):

\[4\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

good, so \[\large \sqrt{50} + \sqrt{32}\] is the same as \[\large 5\sqrt{2} + 4\sqrt{2}\]

OpenStudy (anonymous):

oh,i understand now \[9\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

you're welcome